OpenStudy (anonymous):

21 is 35% of what number ?

OpenStudy (dls):

60

OpenStudy (lgbasallote):

35% is 0.35 21 is 35% of what number let the number = x 21 is 35% of x 21 = 0.35x solve for x

OpenStudy (dls):

35/100*x=21 x=21*100/35=60.

OpenStudy (anonymous):

It's simple just read the first word , 21= 35/100 of x , just solve for x